# GPU-Accelerated Gravitational Lensing &\&& Dynamical (GLaD) Modeling for Cosmology and Galaxies

### v1

@@ 5.1 Breaking the MSD using joint modeling
The BH in lensing-only modeling is often neglected since SL only provides constraints at Einstein radius, which is far from the galaxy’s center. Only in some rare cases, the lensed source image appears close to the galaxy center within $\lesssim 1\leavevmode\nobreak\ \rm kpc$ (e.g., Nightingale et al. 2023; Melo-Carneiro et al. 2025). Kinematic data can provide some constraints, but its effectiveness is highly limited by the instrument’s resolution, particularly for galaxies at galaxies at $z>0.1$. The lens galaxy RXJ1131 at $z=0.295$ might hold a supermassive BH with $M_{\rm BH}$ in the range of [$10^{9},10^{10}]$ ${\rm M_{\odot}}$ which corresponding to a sphere of influence $r_{\rm soi}$ for the BH in the range of $[0.011\arcsec,0.11\arcsec]$. The simulated kinematic data has the spaxel size with $0.1\arcsec$ convolved with $\rm FWHM=0.14\arcsec$ of $\rm PSF_{\rm kin}$. For $M_{\rm BH}$ near $10^{10}\,{\rm M_{\odot}}$, the influence of the BH dynamics can be imprinted on the central Voronoi bins.
@@
As discussed in Section 3.3, we conduct joint modeling for a sequence of $M_{\rm BH}$ values. The time-delay distance $D_{\rm\Delta t,int}$ is almost unaffected by $M_{\rm BH}$. Therefore, we concentrate on the scatter in $D_{\rm d}$ and $\beta_{\rm ani}$, which are constrained exclusively by the kinematic data. In our experiment, the values of $\beta_{\rm ani}$ are distributed across the full prior range of [--0.3, 0.3] (see Fig. 9), given $M_{\rm BH}\in[10^{9},10^{10}]\leavevmode\nobreak\ {\rm M_{\odot}}$. This prior range is motivated by studies of nearby massive elliptical galaxies (see review in Cappellari 2025, figs. 8, 10), and is quite conservative in its broad range compared to the typical scatter of anisotropies of galaxies shown in Cappellari (2025). The anisotropy $\beta_{\rm ani}$ is constrained by the spatial pattern in the kinematic data. However, $M_{\rm BH}$ and $\beta_{\rm ani}$ similarly affect stellar motions in the galaxy centroid, resulting in a trade-off between them. In Fig. 9, we observe that a heavier $M_{\rm BH}$ leads to a smaller $\beta_{\rm ani}$, and vice versa. A higher $M_{\rm BH}$ deepens the central gravitational potential, allowing more tangential orbits in the dynamial model when reproducing the same observed line-of-sight velocity dispersion, corresponding to $\beta_{\rm ani}<0$. Conversely, a lower $M_{\rm BH}$ can produce similar velocity dispersions if the stellar orbits are more radial, with $\beta_{\rm ani}>0$, as radial orbits allow stars to reach higher line-of-sight velocity dispersion near the galaxy center. Both BH mass and $\beta_{\rm ani}$ contribute to accelerating stellar motion, but in different directions.
@@
In addition to its degeneracy with $M_{\mathrm{BH}}$, $\beta_{\mathrm{ani}}$ is also positively correlated with $D_{\rm d}$ (see Fig. 6). An increase in $\beta_{\mathrm{ani}}$ compensates for the influence of a more massive BH by mimicking its effect on stellar motion in the central region. Since we assume a constant $\beta_{\mathrm{ani}}$ across all radii in the joint modeling, this means that even in the outer regions where velocity adjustments are unnecessary, stellar velocities are still affected by $\beta_{\mathrm{ani}}$. To match the observed kinematics beyond the central region, $D_{\rm d}$ increases to counterbalance the effect introduced by changes in $\beta_{\mathrm{ani}}$. This is because $D_{\rm d}$ acts as a normalization factor for scaling $v_{\rm rms}^{\rm pre}$, following the relation $v_{\rm rms}^{\rm pre}\sim\frac{1}{\sqrt{D_{\rm d}}}$ (see Sect. 2.4). If the BH mass in the joint modeling is heavier than the mock input, the entire trend reverses. This explains the observed correlations in the values of $D_{\rm d}$, $\beta_{\rm ani}$ and $M_{\rm BH}$ (see Figs. 4, 5, and 9). In Fig. 5, we observe a negative correlation between the adopted $M_{\rm BH}$ in the joint modeling and $D_{\rm d}$. Black hole masses in the range of $M_{\rm BH}=2\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ to $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ are difficult to distinguish based on kinematic data and all contribute to the inference of distances in the BIC framework, given the $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the mock input.
By combining all joint models weighted by the BIC (i.e., $P(\bm{d}_{\rm D}\mid\mathcal{M}_{i})$, where $\mathcal{M}_{i}$ represents model $i$, either a composite mass model with BH mass $M_{{\rm BH},i}$ or an EPL model), we obtain $D_{\rm d}=781_{-29}^{+30}$ Mpc and $\beta_{\rm ani}=0.21_{-0.14}^{+0.07}$ (see Tab. 4). The recovered distance $D_{\rm d}$ closely matches the input value, whereas the orbital anisotropy $\beta_{\rm ani}$ remains poorly constrained. This is because models with $M_{\rm BH}<5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ tend to cluster near the upper bound of the $\beta_{\rm ani}$ prior (see Fig. 9). Some of these models provide a good fit to the kinematic data and are only slightly downweighted by the BIC, leading to a significant contribution to the final probability density distribution of $\beta_{\rm ani}$.
Figure 9: Marginalized posterior density distribution of $\beta_{\rm ani}$, based on joint models using ideal kinematic data. Different colors represent posterior densities corresponding to different black hole masses, while the same color indicates models with identical mass parameterization but different source grid resolutions. We observe that as $M_{\rm BH}$ increases, the inferred $\beta_{\rm ani}$ decreases and vice versa. The inferred $\beta_{\rm ani}$ distributions given different $M_{\rm BH}$ spread over the prior range [--0.3,0.3], but different $M_{\rm BH}$ yield different goodness of fit to the kinematic data (as shown in Fig. 4).
@@
We investigate the impact of an incorrect BH mass in the joint modeling. First, we test the scenario where we assume no BH in the composite mass model. The inferred value of $H_{0}=83.2_{-3.0}^{+2.3}\rm\,km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$ successfully recovers the input $H_{0}$ value (see Fig. 10 and Table 4). However, the best-fit kinematic map exhibits a significantly different pattern compared to the one obtained when the BH is included in the modeling (see Fig. 7). The difference of $\Delta\chi_{\rm dyn}^{2}=14$ for the dynamical fit indicates a poorer fit compared to the best-fit model that accounts for the BH. The fitted value of $\beta_{\rm ani}=0.29_{-0.006}^{+0.003}$ reaches the upper bound of the prior range, yet it remains insufficient to fully compensate for the absence of the BH, leading to a suboptimal fit to the kinematic data. The high $\beta_{\rm ani}$ value leads to an excessive increase in velocity dispersions in the outer regions. As a result, this imbalance starts to affect the probability density distribution of $D_{\rm\Delta t,int}$. To compensate for the effect induced by $\beta_{\rm ani}$, $D_{\rm\Delta t,int}$ decreases slightly.
@@
Figure 10: Constraints on $H_{0}$ and $\Omega_{\rm m}$ from our models in flat $\Lambda\rm CDM$ cosmology, using the ideal kinematic data and different assumptions for $M_{\rm BH}$. The green contour represents the constraints from the joint modeling, considering the possible values of $M_{\rm BH}$, and combining all 55 models weighted by BIC. The orange contour shows the constraints from the joint modeling excluding the BH mass. The blue contour shows the constraints from the joint modeling with $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$. The BH mass in the simulated kinematics data is set to $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$.
@@
The above tests indicate that a severely misfitted $\beta_{\rm ani}$ can strongly bias $D_{\rm d}$ and mildly influence $D_{\rm\Delta t,int}$ in the extreme case, thereby affecting $H_{0}$ inference, even when the kinematic data appears to be well-fitted. The value of $\beta_{\rm ani}$ can be accurately recovered when the BH mass is known and vice versa. We find that the fitted value of $\beta_{\rm ani}=0.13_{-0.04}^{+0.04}$ is well constrained when the true $M_{\rm BH}$ is used in the joint modeling (see Table 4). However, in nearly all cases of lens galaxies, the precise BH mass is unknown. The bias in $D_{\rm d}$ caused by a misfitted $\beta_{\rm ani}$ can be mitigated by performing joint modeling over a range of possible BH masses and using the BIC to downweight models that are disfavored by the kinematic data. It naturally follows that the prior range of $\beta_{\rm ani}$ should be carefully chosen. Expanding the prior range allows adjustments to $\beta_{\rm ani}$ and $D_{\rm d}$ to always effectively compensate for the presence of the BH. While this results in a well-fitted kinematic model, it significantly biases the inferred $D_{\rm d}$.
We perform joint modeling using the ideal kinematic map while excluding the central regions. We reassess the recovery of the $H_{0}$ value and evaluate the quality of the kinematic fit for both cases of no BH and a BH with $M_{\rm BH}=7\times 10^{9}\,M_{\odot}$. In both cases, we observe that $D_{\rm d}$ and $\beta_{\rm ani}$ shifted closer to the mock input values, allowing for an accurate recovery of $H_{0}$ within $1\sigma$ uncertainties (see Tab. 4 and more details in Appendix. D). As anticipated, the $1\sigma$ uncertainties are broader than the full ideal kinematic dataset because we adopt 43 bins instead of the complete dataset with 52 bins. Additionally, the kinematic data excluding the central region is effectively recovered through joint modeling with no BH and $M_{\rm BH}=7\times 10^{9}\,M_{\odot}$ (see Fig. 11). This suggests that excluding the central kinematic region can help mitigate the effects of the presence of the BH with highly uncertain mass.
As described in Sect. 5.1, we run all modelings using the systematically biased kinematic data. We adopt the composite mass model, $\kappa_{\rm int,comp}$, across the black hole mass range, and a single EPL profile, $\kappa_{\rm int,epl}$, for the joint modeling. To account for degeneracies induced by the source grid resolution, we perform each mass model analysis on source grids ranging from $60\times 60$ to $68\times 68$ pixels. We perform the BIC weighting to combine all 55 joint models. The biased kinematic data helps break the internal MSD, yielding consistent results for $\lambda=0.97_{-0.07}^{+0.04}$ and $D_{\rm\Delta t,int}=1863_{-80}^{+144}$ Mpc, which agree with values inferred from the joint modeling using ideal kinematic data. This demonstrates that the overall systematic bias does not affect the constraints on $D_{\rm\Delta t,int}$ and $\lambda_{\rm int}$ (see Fig. 6). This is because $\lambda_{\rm int}$ is constrained by the 2D kinematic map, where the shape of the $v_{\rm rms}$ profile breaks the internal MSD and constrains $D_{\rm\Delta t,int}$. The 5% bias does not alter the shape of the $v_{\rm rms}$ profile, which is why neither $D_{\rm\Delta t,int}$ nor $\lambda_{\rm int}$ is biased. Following the same reason, the inference of $\beta_{\rm ani}$ remains unaffected. We obtain $\beta_{\rm ani}=0.20_{-0.13}^{+0.08}$, which is consistent with the value inferred using ideal kinematic data.
The systematic bias primarily impacts $D_{\rm d}$ because it changes the amplitude of the $v_{\rm rms}$ overall. Given the relation $v_{\rm rms}^{\rm pre}\sim\frac{1}{\sqrt{D_{\rm d}}}$, a $5\%$ bias in $v_{\rm rms}^{\rm pre}$ results in an expected $\sim 9\%$ bias in $D_{\rm d}$. We obtain $D_{\rm d}=706_{-25}^{+20}$ Mpc, which is $9\%$ lower than the mock input value of $D_{\rm d}=775$ Mpc, as expected. If the combined kinematics are obtained from a single aperture rather than an IFU, the impact on distances will not be cleanly isolated to $D_{\rm d}$ alone, as the single aperture lacks information on the shape of $v_{\rm rms}$. We anticipate a more severe effect on both $D_{\rm d}$ and $D_{\rm\Delta t,int}$
@@ 5.4 The impact of high systematic bias in kinematics data on $H_{0}$ measurement
Table 4: Important parameters and the inferred $H_{0}$ [$\rm kms^{-1}\leavevmode\nobreak\ Mpc^{-1}$] from different joint models. In the individual models, we present the marginalized values of $H_{0}$ constrained by $P(D_{\rm d})$, $P(D_{\rm\Delta t,int})$, and $P(D_{\rm d},D_{\rm\Delta t,int})$, respectively. We also provide the marginalized distance values for $D_{\rm d}$ and $D_{\rm\Delta t,int}$. The $1\sigma$ uncertainties are calculated from the 16th, 50th, and 84th percentiles of the distribution. The input mock values are $H_{0}=82.5\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$, $D_{\rm d}=775$ Mpc and $D_{\rm\Delta t,int}=1823$ Mpc. The star symbol denotes joint modeling that includes the BH mass, which is the mock input.

### v2

@@ 5.3 The impact of high systematic bias in kinematics data on $H_{0}$ measurement
The BH in lensing-only modeling is often neglected since SL only provides constraints at Einstein radius, which is far from the galaxy’s center. Only in some rare cases, the lensed source image appears close to the galaxy center within $\lesssim 1\leavevmode\nobreak\ \rm kpc$ (e.g., Nightingale et al. 2023; Melo-Carneiro et al. 2025). Kinematic data can provide some constraints, but its effectiveness is highly limited by the instrument’s resolution, particularly for galaxies at galaxies at $z>0.1$. The lens galaxy RXJ1131 at $z=0.295$ might hold a supermassive BH with $M_{\rm BH}$ in the range of [$10^{9},10^{10}]$ ${\rm M_{\odot}}$ which corresponding to a sphere of influence $r_{\rm soi}$ for the BH in the range of $[0.011\arcsec,0.11\arcsec]$. The simulated kinematic data has a spaxel size with $0.1\arcsec$ convolved with $\rm FWHM=0.14\arcsec$ of $\rm PSF_{\rm kin}$. For $M_{\rm BH}$ near $10^{10}\,{\rm M_{\odot}}$, the influence of the BH dynamics can be imprinted on the central Voronoi bins.
We investigate the impact of the BH on distance inference using joint modeling with ideal kinematic data. This is due to the well-known mass–anisotropy degeneracy (e.g., Binney & Mamon 1982; Treu & Koopmans 2002; Courteau et al. 2014) in spherical systems, where a wide range of density profiles can be matched by varying $\beta_{\rm ani}$ while keeping the observed $v_{\rm rms}$ unchanged. The lens galaxy in RXJ1131, as a slow rotator with a nearly spherical central structure, is affected by this degeneracy (Cappellari 2016).
@@
The joint modeling results show that the inferred values of $\beta_{\rm ani}$ span the full prior range of [--0.3, 0.3] (see Fig. 8), given a black hole mass in the range $M_{\rm BH}\in[10^{9},10^{10}]{\rm M_{\odot}}$. This prior range is motivated by studies of nearby massive elliptical galaxies (see review in Cappellari 2025, Figs. 8, 10), and is quite conservative in its broad range compared to the typical scatter of anisotropies of galaxies shown in Cappellari (2025). The anisotropy $\beta_{\rm ani}$ is constrained by the spatial pattern in the kinematic data. However, $M_{\rm BH}$ and $\beta_{\rm ani}$ similarly affect stellar motions in the galaxy centroid, resulting in a trade-off between them. In Fig. 8, we observe that a heavier $M_{\rm BH}$ leads to a smaller $\beta_{\rm ani}$, and vice versa. A higher $M_{\rm BH}$ deepens the central gravitational potential, allowing more tangential orbits in the dynamical model when reproducing the same observed line-of-sight velocity dispersion, corresponding to $\beta_{\rm ani}<0$. Conversely, a lower $M_{\rm BH}$ can produce similar velocity dispersions if the stellar orbits are more radial, with $\beta_{\rm ani}>0$, as radial orbits allow stars to reach higher line-of-sight velocity dispersion near the galaxy center. Both BH mass and $\beta_{\rm ani}$ contribute to accelerating stellar motion, but in different directions.
@@
A trade-off between $M_{\rm BH}$ and $\beta_{\rm ani}$ can lead to a misestimated $\beta_{\rm ani}$, which in turn affects the accuracy of the distance inference, particularly in $D_{\rm d}$. When the BH mass used in the modeling is lower than the mock input value, a higher value of $\beta_{\mathrm{ani}}$ is required to fit the kinematic data. Since we assume a constant $\beta_{\mathrm{ani}}$ across all radii in the joint modeling, this means that even in the outer regions where $v_{\rm rms}^{\rm pre}$ adjustments are unnecessary, $v_{\rm rms}^{\rm pre}$ are still affected by $\beta_{\mathrm{ani}}$. To match the observed kinematics beyond the central region, $D_{\rm d}$ increases to counterbalance the effect introduced by changes in $\beta_{\mathrm{ani}}$. This is because $D_{\rm d}$ acts as a normalization factor for scaling $v_{\rm rms}^{\rm pre}$, following the relation $v_{\rm rms}^{\rm pre}\sim\frac{1}{\sqrt{D_{\rm d}}}$ (see Sect. 2.4). If the BH mass in the joint modeling is heavier than the mock input, the entire trend reverses. This explains the observed correlations in the values of $D_{\rm d}$, $\beta_{\rm ani}$ and $M_{\rm BH}$ (see Figs. 5, 8 and 9). In Fig. 9, we observe a negative correlation between the adopted $M_{\rm BH}$ in the joint modeling and $D_{\rm d}$. BH masses in the range of $M_{\rm BH}=2\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ to $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ are difficult to distinguish based on kinematic data and all contribute to the inference of distances in the BIC framework, given the $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$ in the mock input. This results in a slight asymmetry in the probability distribution of $D_{\rm d}$. Since we use $P(\bm{d}_{\rm D}\mid\mathcal{M}_{i})$ for BIC weighting in the joint models, some models with lower or larger BH masses can still provide comparably good fits to the kinematic data as the model with the true $M_{\rm BH}$, by appropriately rescaling both $D_{\rm d}$ and $\beta_{\rm ani}$. These models contribute to the extended tails of the inferred $D_{\rm d}$ distribution (see the upper panels in Fig. 5). Nonetheless, the inferred $D_{\rm d}$ recovers the mock input value within the $1\sigma$ uncertainty. For the same underlying reason—but with a more pronounced impact—the constraint on $\beta_{\rm ani}$ is significantly weaker. As a result, $\beta_{\rm ani}$ is not tightly constrained, with an inferred value of $\beta_{\rm ani}=0.21_{-0.14}^{+0.07}$ (see Tab. 14).
@@
While the above analysis considers a range of plausible BH masses with BIC weighting, we now turn to the case of a single, incorrect BH mass assumption to examine its impact on the inferred distances. First, we test the scenario where we assume no BH in the composite mass model. The inferred value of $H_{0}=83.2_{-3.0}^{+2.3}\rm\,km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$ successfully recovers the input $H_{0}$ value (see Fig. 10 and Table 14). However, the best-fit kinematic map exhibits a significantly different pattern compared to the one obtained when the BH is included in the modeling (see Fig. 6). The difference of $\Delta\chi_{\rm dyn}^{2}=14$ for the dynamical fit indicates a poorer fit compared to the best-fit model that accounts for the BH. The fitted value of $\beta_{\rm ani}=0.29_{-0.006}^{+0.003}$ reaches the upper bound of the prior range, yet it remains insufficient to fully compensate for the absence of the BH, leading to a suboptimal fit to the kinematic data. The high $\beta_{\rm ani}$ value leads to an excessive increase in velocity dispersions in the outer regions. As a result, this imbalance starts to affect the probability density distribution of $D_{\rm\Delta t,int}$. To compensate for the effect induced by $\beta_{\rm ani}$, $D_{\rm\Delta t,int}$ decreases slightly.
@@
Figure 10: Constraints on $H_{0}$ and $\Omega_{\rm m}$ from our models in flat $\Lambda\rm CDM$ cosmology, using the ideal kinematic data and different assumptions for $M_{\rm BH}$. The green contour represents the constraints from the joint modeling, considering the possible values of $M_{\rm BH}$, and combining all 55 models weighted by BIC. The orange contour shows the constraints from the joint modeling, excluding the BH mass. The blue contour shows the constraints from the joint modeling with $M_{\rm BH}=7\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$. The BH mass in the simulated kinematics data is set to $M_{\rm BH}=5\times 10^{9}\leavevmode\nobreak\ {\rm M_{\odot}}$.
@@
The above tests indicate that a severely misfitted $\beta_{\rm ani}$ can strongly bias $D_{\rm d}$ and mildly influence $D_{\rm\Delta t,int}$ in the extreme case, thereby affecting $H_{0}$ inference, even when the kinematic data appears to be well-fitted. The value of $\beta_{\rm ani}$ can be accurately recovered when the BH mass is known and vice versa. We find that the fitted value of $\beta_{\rm ani}=0.13_{-0.04}^{+0.04}$ is well constrained when the true $M_{\rm BH}$ is used in the joint modeling (see Table 14). However, in nearly all cases of lens galaxies, the precise BH mass is unknown. The bias in $D_{\rm d}$ caused by a misfitted $\beta_{\rm ani}$ can be mitigated by performing joint modeling over a range of possible BH masses and using the BIC to downweight models that are disfavored by the kinematic data. It naturally follows that the prior range of $\beta_{\rm ani}$ should be carefully chosen Simon et al. (2024). Expanding the prior range allows adjustments to $\beta_{\rm ani}$ and $D_{\rm d}$ to always effectively compensate for the presence of the BH. While this results in a well-fitted kinematic model, it significantly biases the inferred $D_{\rm d}$.
Table 4: Inferred $H_{0}$ and key parameters across models.
@@
 [^14] Important parameters and the inferred $H_{0}$ [$\rm kms^{-1}\leavevmode\nobreak\ Mpc^{-1}$] from different joint models. In the individual models, we present the marginalized values of $H_{0}$ constrained by $P(D_{\rm d})$, $P(D_{\rm\Delta t,int})$, and $P(D_{\rm d},D_{\rm\Delta t,int})$, respectively. We also provide the marginalized distance values for $D_{\rm d}$ and $D_{\rm\Delta t,int}$. The $1\sigma$ uncertainties are calculated from the 16th, 50th, and 84th percentiles of the distribution. The input mock values are $H_{0}=82.5\leavevmode\nobreak\ \rm km\leavevmode\nobreak\ s^{-1}\leavevmode\nobreak\ Mpc^{-1}$, $D_{\rm d}=775$ Mpc and $D_{\rm\Delta t,int}=1823$ Mpc. The star symbol denotes joint modeling that includes the BH mass, which is the mock input.
We perform joint modeling using the ideal kinematic map while excluding the central regions. We reassess the recovery of the $H_{0}$ value and evaluate the quality of the kinematic fit for both cases of no BH and a BH with $M_{\rm BH}=7\times 10^{9}\,M_{\odot}$. In both cases, we observe that $D_{\rm d}$ and $\beta_{\rm ani}$ shifted closer to the mock input values, allowing for an accurate recovery of $H_{0}$ within $1\sigma$ uncertainties (see Tab. 14 and more details in Appendix. E). As anticipated, the $1\sigma$ uncertainties are broader than the full ideal kinematic dataset because we adopt 43 bins instead of the complete dataset with 52 bins. Additionally, the kinematic data excluding the central region is effectively recovered through joint modeling with no BH and $M_{\rm BH}=7\times 10^{9}\,M_{\odot}$ (see Fig. 11). This suggests that excluding the central kinematic region can help mitigate the impact of a BH with a highly uncertain mass, and reduce potential biases in distance inference. However, in both cases, $\beta_{\rm ani}$ remains poorly constrained. This is due to the fact that the outer regions of galaxies are often significantly influenced by dark matter. The presence of dark matter introduces its own mass–anisotropy degeneracy, which complicates the interpretation of kinematic data in these regions—similar to the effect of the BH in the central regions. As a result, determining $\beta_{\rm ani}$ using only the outer kinematic bins is challenging due to the unknown distribution of dark matter.
To fully resolve the $M_{\rm BH}$$\beta_{\rm ani}$ degeneracy, one effective way is to use the full line-of-sight velocity distribution (LOSVD). Instead of relying only on the second velocity moments $v_{\rm rms}$, incorporating higher-order moments like $h_{3}$ and $h_{4}$ provides stronger constraints on the orbital structure. This approach, however, demands high signal-to-noise data and advanced modeling techniques, such as Schwarzschild’s orbit-superposition method (Cappellari et al. 2009; Thomas et al. 2014, 2016; Poci & Smith 2022). The lens galaxy of RXJ1131 at $z=0.295$ even with JWST NIRSpec data is unlikely to reach the required resolution and signal-to-noise ratio to perform such modeling.
Another approach, while still relying only on $v_{\rm rms}$ in JAM modeling, is to adopt a physically motivated prior for a more flexible orbital anisotropy. This can be informed by studies of well-observed nearby galaxies, as demonstrated in Simon et al. (2024). In that work, the author introduced a radially varying anisotropy profile, $\beta_{\rm ani}(r)$, and carefully defined the prior range based on physical expectations. Specifically, the central regions near the black hole are assumed to be more tangentially biased, while the outer regions are expected to be more radially anisotropic. This pattern is supported by theoretical predictions that dry mergers eject stars on radial orbits, which pass close to a central supermassive black hole binary. Such interactions can create a core in the surface brightness profile and lead to a dominance of tangential orbits in the central region, while the stars in the outer regions retain a more radial orbital distribution. This scenario is consistent with both observations and simulations (e.g. Milosavljević & Merritt 2001; Rantala et al. 2018; Saglia et al. 2024).
In principle, a similar methodology, incorporating well-justified anisotropy priors, could also be useful in lensing-based dynamical analyses. However, the current version of our modeling code supports only a constant anisotropy across radii. Fully resolving the $M_{\rm BH}$$\beta_{\rm ani}$ degeneracy is beyond the scope of this paper. Instead, the main goal of this work is to show that, once this degeneracy is reasonably mitigated, it does not introduce significant bias on the inferred cosmological distances.